OpenStudy (anonymous):

2sin^2x-cosx-1=0 my answer have to be redians

OpenStudy (jdoe0001):

\(\bf 2sin^2(x)-cos(x)-1=0\qquad \textit{recall that}\quad sin^2(\theta)=1-cos^2(\theta)\\ \quad \\ 2sin^2(x)-cos(x)-1=0\implies 2[1-cos^2(x)]-cos(x)-1=0\\ \quad \\ -2cos^2(x)-cos(x)-1+2=0\) tis a quadratic, so factor it likewise